OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 3y = 8? @rm152

OpenStudy (confusionist):

I can't do this unless I can see the graphs. Can you link them? :)

OpenStudy (anonymous):

i would have to saved 4 photos can u graph and ill figure it out? @Confusionist

OpenStudy (confusionist):

You want me to graph 3y=8 for you? :o I'm confused. If you upload the images, I can do that, too.

OpenStudy (anonymous):

there are four graphs it would be a long process can you just draw it can't you do that? @Confusionist

OpenStudy (confusionist):

Sorry, I'm on a touchscreen laptop. It doesn't draw. Hopefully someone else can help you! :)

OpenStudy (anonymous):

ill post two and you see if they match

OpenStudy (anonymous):

look at this

OpenStudy (anonymous):

@Confusionist are you helping?

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (confusionist):

I said I wasn't able to help. Sorry! Hopefully someone else can. :)

OpenStudy (inkyvoyd):

Mind posting the original problem? If the problem gives 4 pictures, please show them all, otherwise I will have trouble helping.

OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 3^y = 8?

Directrix (directrix):

@woohoo Here is a graph of 3^y = 8. Is that what you are asking for? http://www.wolframalpha.com/input/?i=+3%5Ey+%3D+8

OpenStudy (confusionist):

The answer is neither of those. ^

OpenStudy (anonymous):

@Directrix that doesnt look like any of the options

OpenStudy (inkyvoyd):

so @woohoo you are basically looking for a base 3 log graph

OpenStudy (inkyvoyd):

take the base 3 log on both sides of 3^y=8 to get y=base 3 log 8

OpenStudy (inkyvoyd):

to find the curve that is base three log 8 you should look at some points; for instance base 3 log 9=2, and base 3 log 3=1

OpenStudy (inkyvoyd):

and base 3 log 1=0

OpenStudy (confusionist):

Rather than just eyeballing it, look at the points on the graph. Like @inkyvoyd said, you'll eventually see familiar points that we've mentioned. Don't look for exact replicas.

OpenStudy (inkyvoyd):

So you should get a graph with points (1,0) (3,1) (9,2)

OpenStudy (inkyvoyd):

I can tell you right now that 8 is a lot closer to 9 than it is 3, so you should see a point (8,approx. 2)

OpenStudy (anonymous):

yeah it looks like 8,2

OpenStudy (anonymous):

so its the ssecond last option

OpenStudy (inkyvoyd):

I would say it is.